### v2

@@ B.1 Detail of Problem Scraping
B.2 Detail of Solution Scraping
During the problem-scraping process, we also collected the links under the “Contest material” section on the right side of each problem page and identified which of those led to solution blogs. Crawling these editorial pages is fairly complex, because Codeforces’ official write-ups are hosted as personal blogs whose formats can vary over time, making content extraction more challenging. Although each round’s problems (e.g. “123A,” “123B,” etc.) live on separate pages, all of the editorials for a given numeric ID usually appear on a single blog page. We therefore need to assign each sub-problem’s write-up (A, B, C, …) to its corresponding problem. To do this, our crawler first locates the distinct anchor points that mark each sub-problem section, then extracts the content between each anchor and the next as that problem’s editorial. The benefit of this approach is that, while in the end we only need the solution code itself, editorials often consist of plain text explanations, a single code snippet, or multiple code variants and languages. Relying on a purely mechanical scraper makes it difficult to isolate exactly the code we want, so it’s more effective to pass the raw editorial content to an LLM for final organization and extraction.
@@ B.2 Detail of Solution Scraping
Therefore, our primary and most precise method is to search for a URL containing the problem ID (for example, “problemset/problem/2060/A” or “contest/blog/2060/A”) and treat its position as the anchor. Once the anchor is found, we scan the surrounding page and extract its content to obtain the editorial.Anchor scanning and recording also relies on a problem-ID reference table, which we built from the IDs of all problems scraped in the first step. Its main role is to guide the code when matching anchors: for instance, if the table shows that numeric ID 2060 has sub-IDs “A” through “G”, the scraper first reads those sub-IDs, then walks through the page using the URL-based or text-based method to record the exact anchor for each sub-ID.An example that fit URL anchor identification technique in shown in Figure 8(a).
However, there are still little parts of early editorial blogs didn’t include URL hyperlinks, so in those cases we fall back on regular-expression–based text matching wherever possible. Based on the formats we observed, we designed two main matching strategies:
  • Difficulty-label matching. A number of blogs publish all of a round’s problems on one page and mark them with labels like “Div2” or “Div1” (since most rounds contain two Div2-level problems and several Div1-level ones). To handle this, we use our problem-ID reference table to identify all sub-IDs belonging to the same numeric contest but with different difficulty levels, tag them accordingly in the table, and then, during the anchor-matching process, if the scraper detects a difficulty label it will also try to match anchors based on that label. An example is shown in Figure 8(b).
  • Problem-label matching. Beyond difficulty tags, many blogs use the literal “Problem A”, “Problem B”, etc. format. We include a specific regex pattern to detect those “Problem+sub-ID” labels and assign each section to the correct sub-problem.
@@
B.3 Detail of Solution Code Generation
@@
B.4 Detail of Subproblems Generation
@@
B.5 Detail of Test Cases Generation
@@
Appendix C Mathematical Expression of $PD@k$& $APD@k$
@@
The probability that $M=d_{(j)}$ equals the probability of choosing $d_{(j)}$ along with $k-1$ depths from the first $j-1$ smaller values:
@@
Appendix D Experiment
D.1 The Setting of Experiment
@@ D.2 Instruction Templates
D.3 Pass Turn Result
Table 6: Multi-turn pass@turn results for various models.
Model Average Turn-1 Turn-2 Turn-3 Turn-4 Turn-5
Closed-Source
o3-mini 0.600 0.322 0.632 0.777 0.983 0.571
o1-mini 0.581 0.233 0.645 0.798 0.879 0.667
GPT-4.1-mini 0.646 0.265 0.760 0.803 1.034 0.762
GPT-4o 0.537 0.177 0.604 0.697 1.017 0.476
GPT-4o-mini 0.467 0.137 0.501 0.638 0.982 0.476
Gemini-2.0-flash 0.541 0.183 0.595 0.702 0.948 0.667
Claude-3.5-Sonnet 0.461 0.117 0.508 0.574 1.000 0.571
Open-Source (7B-Level)
Qwen2.5-Coder-7B-Instruct 0.258 0.018 0.270 0.394 0.638 0.286
Llama-3.1-8B-Instruct 0.232 0.011 0.245 0.404 0.534 0.476
Yi-Coder-9B-Chat 0.247 0.021 0.314 0.324 0.362 0.286
Open-Source (32B-Level)
Qwen2.5-Coder-32B-Instruct 0.352 0.067 0.391 0.532 0.569 0.666
QwQ-32B-Preview 0.515 0.261 0.606 0.553 0.828 0.381
Open-Source (70B-Level)
Llama-3.3-70B-Instruct 0.493 0.163 0.515 0.681 1.000 0.476
Qwen2.5-72B-Instruct 0.330 0.110 0.350 0.452 0.517 0.429
Open-Source (Large Model)
Deepseek-V3 0.572 0.219 0.622 0.750 0.966 0.810
Deepseek-R1 0.609 0.304 0.677 0.766 0.966 0.860
Appendix E Discussion
E.1 Discussion on Models’ DSC Performance
@@
[…]
@@
E.2 Error Case Study
@@
Appendix F Impact Statement
@@
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The main claims are clearly presented in the abstract and introduction, and are further elaborated and substantiated throughout the subsequent sections of the paper.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: Please refer to the Section 5 in the paper.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: The paper does not include theoretical results.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: Please refer to the Section 3.1 and Appendix D.1 in the paper.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We provide well-documented repositories to store dataset and code. The URL can be found in the first page.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: Please refer to the Section 4 and Appendix D.1, D.2 in the paper.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: Please refer to the Section 4 and Appendix D.1 in the paper.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification:Please refer to the Section 4 and Appendix D.1 in the paper.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: This paper conforms, in every respect, with the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: Please refer to the Appendix F in the paper.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: The paper poses no such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: The creators and original owners of all assets used in the paper are properly credited, and the licenses and terms of use are explicitly mentioned and fully respected.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: The paper does not release new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper doesn’t include crowdsourcing or research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [Yes]
  79. Justification: LLM is used as a tool in data curation pipeline. Please refer to the Section 3.1 and Appendix B.3, B.4 in the paper.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.

### v3

@@ A.1 Detail of Problem Scraping
A.2 Detail of Solution Scraping
During the problem‐scraping process, we also collected the links under the “Contest material” section on the right side of each problem page and identified which of those led to solution blogs. Crawling these editorial pages is fairly complex, because Codeforces’ official write‐ups are hosted as personal blogs whose formats can vary over time, making content extraction more challenging. Although each round’s problems (e.g. “123A,” “123B,” etc.) live on separate pages, all of the editorials for a given numeric ID usually appear on a single blog page. We therefore need to assign each sub‐problem’s write‐up (A, B, C, …) to its corresponding problem. To do this, our crawler first locates the distinct anchor points that mark each sub‐problem section, then extracts the content between each anchor and the next as that problem’s editorial. The benefit of this approach is that, while in the end we only need the solution code itself, editorials often consist of plain text explanations, a single code snippet, or multiple code variants and languages. Relying on a purely mechanical scraper makes it difficult to isolate exactly the code we want, so it’s more effective to pass the raw editorial content to an LLM for final organization and extraction.
@@ A.2 Detail of Solution Scraping
Therefore, our primary and most precise method is to search for a URL containing the problem ID (for example, “problemset/problem/2060/A” or “contest/blog/2060/A”) and treat its position as the anchor. Once the anchor is found, we scan the surrounding page and extract its content to obtain the editorial.Anchor scanning and recording also relies on a problem‑ID reference table, which we built from the IDs of all problems scraped in the first step. Its main role is to guide the code when matching anchors: for instance, if the table shows that numeric ID 2060 has sub‑IDs “A” through “G”, the scraper first reads those sub‑IDs, then walks through the page using the URL‑based or text‑based method to record the exact anchor for each sub‑ID.An example that fit URL anchor identification technique in shown in Figure 7(a).
However, there are still little parts of early editorial blogs didn’t include URL hyperlinks, so in those cases we fall back on regular‑expression–based text matching wherever possible. Based on the formats we observed, we designed two main matching strategies:
  • Difficulty‑label matching. A number of blogs publish all of a round’s problems on one page and mark them with labels like “Div2” or “Div1” (since most rounds contain two Div2‑level problems and several Div1‑level ones). To handle this, we use our problem‑ID reference table to identify all sub‑IDs belonging to the same numeric contest but with different difficulty levels, tag them accordingly in the table, and then, during the anchor‑matching process, if the scraper detects a difficulty label it will also try to match anchors based on that label. An example is shown in Figure 7(b).
  • Problem‑label matching. Beyond difficulty tags, many blogs use the literal “Problem A”, “Problem B”, etc. format. We include a specific regex pattern to detect those “Problem+sub‑ID” labels and assign each section to the correct sub‑problem.
@@
A.3 Detail of Solution Code Generation
@@
A.4 Detail of Subproblems Generation
@@
A.5 Detail of Test Cases Generation
@@
A.6 Ablation Study of LLM usage
First, to check its influence on the content and style of Subproblem Descriptions, we compared Deepseek-V3 against GPT-4o-mini, Gemini-2.0-Flash and o3-mini. As shown in Table 4, the high semantic overlap and consistent stylistic metrics indicate that descriptions are primarily driven by the underlying code logic rather than model-specific habits. The main content of subproblems is not highly influenced by model choice.
Second, to check its influence on the structure of LLM-generated solution code, we re-generated solutions using o3-mini, Claude-3.7-Sonnet, and Gemini-2.5-Pro. The resulting complexity metrics (Table 5) were identical across all models, confirming that the model produces almost identical decomposition results, indicating very minor bias between different LLMs.
Finally, to verify that our subproblems are natural and representative of human problem-solving, we conducted a Human-LLM cross-validation by constructing a creteria including five aspects. As shown in Table 6, both human and model judges consistently rated the subproblems as high-quality, which confirms that our automated pipeline produces professional, human-aligned problem decompositions.
Prompts of Subproblem Description Quality Validation is shown as follows:
Subproblem Description Quality Validation You are an expert evaluator of programming task subproblems. You will receive a subproblem (including the statement, solution code, and test cases). Evaluate whether the statement adequately meets each of the following five criteria. For each criterion, assign **1** if it is satisfied, or **0** if it is not. The total possible score is thus 5. 1. Clarity: Is the statement smoothly and understandably expressed, allowing a reader to grasp the task goal quickly? 2. Completeness: Does the description include all key elements needed to accomplish the subtask? 3. Accuracy: Is the description free of ambiguity or logical errors, and does it match the problem requirements and the provided code? 4. Feasibility: Can an engineer unambiguously determine and implement the required functionality—and pass functional tests—based solely on this description? 5. Professionalism: Does it use accurate, domain-appropriate terminology and a style fitting technical norms of coding tasks? Return your result **only** as a JSON dictionary with these five keys and values of 0 or 1. Please do not be overly strict. Assign a score of 1 to a criterion if it is even partially satisfied, allowing for minor imperfections. Only give a score of 0 if there are major or fundamental issues. Furthermore, ensure that the overall score for any subproblem is not lower than 2. Here is the subproblem:content Only return the JSON dictionary and remove the “‘json note, nothing else.
Table 4: Robustness of Subproblem Descriptions (Ref: Deepseek-V3). Left: Semantic Consistency; Right: Stylistic Similarity.
Model Semantic Stylistic
ROUGE-1 BERT Len. Ratio Marker Diff. Error Diff.
GPT-4o-mini 0.582 0.725 0.975 0.026 0.367
Gemini-2.0-Flash 0.600 0.734 1.141 0.019 0.728
o3-mini 0.554 0.692 1.181 0.124 1.350
Table 5: Structural Complexity Metrics Across Models
Model Avg. Turns Avg. Depth
o3-mini 2.08 1.94
Claude-3.7-Sonnet 2.08 1.94
Gemini-2.5-Pro 2.08 1.94
Table 6: Human Evaluation of Subproblem Quality. Metrics include Clarity (Cla.), Completeness (Com.), Accuracy (Acc.), Feasibility (Fea.), and Professionalism (Pro.).
Source Avg. Rating Cla. Com. Acc. Fea. Pro.
o3-mini 4.222 0.981 0.604 0.955 0.797 0.986
Claude-3.7 4.038 0.889 0.745 0.745 0.668 0.990
Human 4.202 0.981 0.659 0.904 0.726 0.933
A.7 Compute Cost Estimation
The cost of using LLM in our data curation pipeline is estimated as below:
  • Generating Solution Code: On average, each problem requires approximately 3K input tokens and 5K output tokens (including reasoning). The official Deepseek-R1 API is currently 4 CNY per 1M input tokens and 16 CNY per 1M output tokens. Therefore, the total cost for 5K+ problems is approximately 450 CNY.
  • Generating Subproblem Description: On average, each problem requires approximately 3K input tokens and 1K output tokens. The official Deepseek-V3 API is currently 2 CNY per 1M input tokens and 8 CNY per 1M output tokens. Therefore, the total cost for 5K+ problems is approximately 70 CNY.
The cost of processing the initial batch of 5K+ problems is manageable and also a one-time effort that does not need to be repeated. Future updates to the benchmark will involve only newly published problems and will therefore be much smaller in scale and less resource-intensive.
Appendix B CodeFlowBench-Comp Dataset
Fig. 12 demonstrates the statistical information of CodeFlowBench-main. Regarding the distribution, the data shows a clear preference for multi-step logic. The overall-turns histogram peaks at 2 turns (accounting for $\sim$2,100 problems), significantly outnumbering single-turn instances ($\sim$1,500), and exhibits a long-tail pattern extending to $\geq$7 turns. Similarly, the overall-depth distribution is heavily concentrated at Depth 2 (>>2,700 problems), indicating that nested dependency chains are the norm.
Furthermore, correlating these metrics with Codeforces difficulty ratings confirms their validity. As illustrated in the rating curves, we observe a strong positive correlation beyond the structural baseline: as the interaction turns and dependency depth increase, the average problem rating rises sharply—approaching 2,900 for tasks with $\geq$7 turns. This demonstrates that our structural metrics effectively capture the intrinsic complexity of logical problems.
(a) Distributions of overall-turns and overall-depth.
(b) The Correlations with Rating Levels.
Figure 12: Statistics of the overall-turns and overall-depth metrics in CodeFlowBench-Comp. Subfigure (b) shows inflection points at turns = 1 and depth = 1. This is attributed to the fact that competition-level problems are not restricted to multi-turn or deeply nested structures. Partial difficult problems are designed to be solvable by single function such as number theory related problems.
Appendix C CodeFlowBench-Repo Dataset
In this section, we provide a detailed overview of the CodeFlowBench-Repo dataset.
C.1 DomainEval
Overview of DomainEval
DomainEval is an auto-constructed benchmark specifically designed to evaluate LLMs’ coding capabilities across diverse real-world domains. Unlike general-purpose benchmarks that focus on common algorithmic tasks, DomainEval targets six distinct domains: Computation, Network, Basic, System, Visualization, and Cryptography. The dataset is curated from representative open-source GitHub repositories that possess high community recognition (typically over 100 stars) to ensure code quality.
Structurally, each subject in DomainEval consists of three essential components: a natural language instruction, a reference solution (target function), and a suite of valid test cases extracted via a test-method matching strategy. Crucially, to ensure executability, DomainEval preserves the necessary context required to run the target code. This design ensures that the tasks are not isolated snippets but are deeply embedded in the dependency structures of authentic software projects.
Rationale for Source Selection.
The selection of DomainEval as the foundation for CodeFlowBench-Domain is a deliberate design choice driven by two key factors:
1. Authenticity across Specialized Domains. First, DomainEval provides a rigorous source of ecologically valid code samples. By extracting problems from mature, production-level repositories, it captures the authentic coding patterns that professional developers actually employ. Furthermore, its explicit coverage of specialized domains like Cryptography and System areas where general LLMs often struggle allows our benchmark to assess Dependency Awareness in contexts that require precise API knowledge, effectively complementing the algorithmic logic focus of standard competition datasets.
2. Alignment with Constructive Codeflow Paradigm.
While many repository-level benchmarks focus on maintenance tasks such as issue resolution or debugging within existing codebases, CodeFlowBench aims to evaluate the constructive aspect of software engineering, namely to build complex functionality from the ground up. DomainEval focuses on function generation within a rich context. This format is the ideal input for our pipeline, allowing us to decompose a monolithic, dependency-heavy function into a step-by-step, iterative construction process. By leveraging DomainEval, we effectively demonstrate that the iterative codeflow paradigm is adaptable to building functional modules in authentic software libraries, validating our pipeline’s effectiveness beyond algorithmic puzzles.
C.2 Dataset Construction
The construction of the CodeFlowBench-Repo subset adapts the Subproblem Generation Phase from the CodeFlowBench-Comp pipeline. A example of problems in DomainEval is shown in Fig. 13
Figure 13: An example problem in DomainEval. We use the "method_code" attribute to generate subproblem solution code and use test case in "test_code" to generate subproblems’ test cases.
Since DomainEval already provides high-quality reference functions as ground truth, we omit the verification step and proceed directly to dependency analysis. We remove problems which only include single-turn function from the dataset and for the remaining, we parse the dependency relationship of the reference code to identify internal function calls and decompose them into iterative subproblems. Similarly, we then call LLM to generate problem description for subproblems. Finally, we use test cases of the top-level function to generate test cases of subproblems. An example is shown in Fig. 14.
Figure 14: An example problem in CodeFlowBench-Repo, which has a shared form of problems in CodeFlowBench-Repo.
To ensure robust evaluation, we parse the import statements from the original test files and utilize Docker containers to encapsulate the complex package dependencies required by these real-world repositories. This ensures a compatible and reproducible execution environment for all generated subproblems.
C.3 Dataset Statistics
Since our pipeline explicitly looks for internal function reuse to create multi-turn steps, the API-centric nature of the Computation domain resulted in empty dependency trees. Consequently, our selection was restricted to the remaining 749 problems from the other five domains. After further filtering, only 70 high-quality multi-turn problems survived the pipeline. To ensure a balanced evaluation, we augmented this set by add 12 single-turn problems each domain and curate the final CodeFlowBench-Repo.
As illustrated in Fig. 15, the dataset features a balanced distribution across five distinct domains: System (28.7%) and Cryptography (23.3%) constitute the majority, followed by Network (21.7%), Basic utilities (17.8%), and Visualization (8.5%). This distribution ensures the benchmark covers a wide spectrum of software engineering scenarios, from low-level OS interactions to high-level data plotting.
Structurally, the dataset presents significant complexity beyond atomic code generation. The distribution of overall turns and overall depths is shown in Fig. 15. The distribution of overall turns reveals that 56.6% of the problems require multi-turn iterative generation, forcing models to maintain reasoning continuity over long horizons. Furthermore, the dependency depth distribution confirms the intricate nature of these tasks: while 45.7% of problems act as foundational nodes (Depth 1), the remaining 54.3% involve deeper dependency chains (Depth 2–4), requiring the model to correctly implement and call prerequisite functions within a hierarchical context.
Figure 15: Statistics of CodeFlowBench-Repo. Left: The diversity of domains, with a focus on System and Cryptography tasks. Right: Distributions of interaction turns and dependency depths, highlighting the structural complexity.
Appendix D Mathematical Expression of PD@kPD@k& $APD@k$
@@
The probability that $M=d_{(j)}$ equals the probability of choosing $d_{(j)}$ along with k1k-1 depths from the first j1j-1 smaller values:
@@
Appendix E Experiment
E.1 Experiment Setting
@@ E.2 Instruction Templates
Appendix F Discussion
F.1 Discussion on Models’ DSC Performance
@@
$\text{DSC(problem)}=\frac{\text{Overall-Turns(problem)}}{\text{Overall-Depth(problem)}}$
@@
F.2 Error Case Study
@@
Appendix G Impact Statement